OpenStudy (anonymous):

New question guys last one of the night....yes you wanna do it....just go on ahead and do it...you know you wanna..... A square poster has sides measuring 2 feet less than the sides of a square sign. If the difference between their areas is 24sq feet. find the lengths of the sides and the posters of the signs.

OpenStudy (anonymous):

let x be the sides of the poster, so area of poster is x^2 then x+2 is the length of the sides of the sign, so area of sign is (x+2)^2 the different in their areas is (x+2)^2 - x^2 = x^2 + 4x + 4 - x^2 = 4x+4 so 4x+4 = 24, solve for x